OpenStudy (condor34):

Find the sum of a 10-term geometric sequence when the first term is 3 and the last term is 59,049 and select the correct answer below. answer choices: 177,147 88,572 88,575 177,144

OpenStudy (anonymous):

an=a1⋅rn−1

OpenStudy (anonymous):

an=3⋅rn−1

OpenStudy (anonymous):

a10=3⋅r9

OpenStudy (anonymous):

59049=3⋅r9

OpenStudy (anonymous):

59049/3=r9

OpenStudy (anonymous):

19683=r9

OpenStudy (anonymous):

r=3

OpenStudy (anonymous):

S10=3(1−310)1−3=88572

OpenStudy (condor34):

can you help me with another?

OpenStudy (anonymous):

sure

OpenStudy (condor34):

Evaluate S5 for 400 + 200 + 100 + … and select the correct answer below. answer choices: 25 775 1,125 500

OpenStudy (anonymous):

so the rate is 1/2 every time

OpenStudy (condor34):

and that is s sub 5 or 2_5

OpenStudy (anonymous):

so the last two numbers are 50 and 25

OpenStudy (anonymous):

so 400 + 200 + 100 + 50+25

OpenStudy (anonymous):

775

OpenStudy (condor34):

thanks

OpenStudy (anonymous):

yea np